OpenStudy (anonymous):

3(x + 5) - 2(x + 2) = 20 solve for x

OpenStudy (anonymous):

Simplifying 3(x + 5) + -2(x + 2) = 20 Reorder the terms: 3(5 + x) + -2(x + 2) = 20 (5 * 3 + x * 3) + -2(x + 2) = 20 (15 + 3x) + -2(x + 2) = 20 Reorder the terms: 15 + 3x + -2(2 + x) = 20 15 + 3x + (2 * -2 + x * -2) = 20 15 + 3x + (-4 + -2x) = 20 Reorder the terms: 15 + -4 + 3x + -2x = 20 Combine like terms: 15 + -4 = 11 11 + 3x + -2x = 20 Combine like terms: 3x + -2x = 1x 11 + 1x = 20 Solving 11 + 1x = 20 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-11' to each side of the equation. 11 + -11 + 1x = 20 + -11 Combine like terms: 11 + -11 = 0 0 + 1x = 20 + -11 1x = 20 + -11 Combine like terms: 20 + -11 = 9 1x = 9 Divide each side by '1'. x = 9 Simplifying x = 9

OpenStudy (stuck-help):